OpenStudy (alfie):

I'm failing with polynomials integrals >.<' No substitution or parts allowed. \[\int {\frac{{3x + 1}}{{{x^2} - 4x + 3}}dx} \] I don't need the result, I'd like someone to explain the reasoning. Thanks.

OpenStudy (lalaly):

partial fractions

OpenStudy (lalaly):

\[\frac{3x+1}{x^2-4x+3} = \frac{3x+1}{(x-3)(x-1)} = \frac{A}{x-3}+\frac{B}{x-1}\]

OpenStudy (lalaly):

do u know how to find A and B?

OpenStudy (alfie):

Yes, I know. I actually knew how to do this method as well, and I thank you for helping out. Just one thing, my book does it in a different way, it actually wants to operate at the numerator and the denominator, I understand the steps it does, but I can't understand where it gets it from (intuition maybe?) going to write them down, sec.

OpenStudy (anonymous):

Perhaps they are using the heaviside method

OpenStudy (lalaly):

ok write them down:)

OpenStudy (alfie):

\[\int {\frac{{3x + 1}}{{{x^2} - 4x + 3}}dx} = \frac{3}{2}\int \frac{{(2x - 4) + \frac{2}{3} + 4}}{{{x^2} - 4x + 3}}dx = \frac{3}{2}\int \frac{{2x - 4}}{{{x^2} - 4x + 3}}dx + 7\int \frac{1}{{{x^2} - 4x + 3}}\]

OpenStudy (alfie):

there's a +7 integral of 1 over same denominator. Then it claims Discriminant is > 0 and does other steps, sec.

OpenStudy (alfie):

\[\begin{gathered} \int {\frac{{3x + 1}}{{{x^2} - 4x + 3}}dx} = \frac{3}{2}\smallint \frac{{(2x - 4) + \frac{2}{3} + 4}}{{{x^2} - 4x + 3}}dx = \\ \frac{3}{2}\smallint \frac{{2x - 4}}{{{x^2} - 4x + 3}}dx + 7\smallint \frac{1}{{{x^2} - 4x + 3}} \smallint \frac{1}{{{x^2} - 4x + 3}}dx = \\ \smallint \frac{1}{{{{(x - 2)}^2} - 1}}dx = - sett - tgh(x - 2) + c \ = \\ \frac{3}{2}\log |{x^2} - 4x + 3| - 7sett,tgh(x - 2) + c \end{gathered} \]

OpenStudy (amistre64):

oy thats an eyesore

OpenStudy (alfie):

and my book result doesn't even match with wolframalpha's one... true story. More over, I have no idea what is this "sett" thingy.

OpenStudy (alfie):

I'll just take it as my book is wrong or something? I am kind of confused on what to do >.<'

OpenStudy (amistre64):

no typos?

OpenStudy (alfie):

I'll take a scan of the page.

OpenStudy (amistre64):

\begin{array}l \frac{3}{2}\int \frac{{(2x - 4) + \frac{2}{3} + 4}}{{{x^2} - 4x + 3}}\\\\ \frac{3}{2}\int \frac{{(2x - 4) + \frac{14}{3}}}{{{x^2} - 4x + 3}}\\\\ \frac{3}{2}\int \frac{{\frac{14}{3}}}{{{x^2} - 4x + 3}}\\\\ \frac{3}{2}\frac{14}{3}\int \frac{{1}}{{{x^2} - 4x + 3}}\\\\ 7\int \frac{{1}}{{{x^2} - 4x + 3}}\\ \end{array} i got it to this far

OpenStudy (amistre64):

i got no idea whay you double up the integral after that

OpenStudy (alfie):

here it is. by the way thank you guys

OpenStudy (mr.math):

I see what your book is trying to do. Your book is trying to make the numerator as the derivative of the denominator plus some constant. That's why it divides the whole thing by \(\frac{3}{2}\). So they can have 2x-4, (which is the derivative of the bottom), plus some number. The way lalaly used is better anyways, esp in the case we have here.

OpenStudy (amistre64):

tg h: tangent hyperbolic?

OpenStudy (mr.math):

Inverse tanh.

OpenStudy (mr.math):

What notations on earth are your book using?

OpenStudy (alfie):

I don't understand, is that possible that there is some book epic failure ? Because this is so weird, more over I don't get why the result is different from wolfram.

OpenStudy (mr.math):

I have to go eat.

OpenStudy (amistre64):

they turned it into a log for the first part to make life simpler; the last part is this \[\int \frac{{1}}{{{x^2} - 4x + 3}}\] complete the square \[\int \frac{{1}}{x^2 - 4x+4-4 + 3}\] \[\int \frac{{1}}{(x-2)^2-1}dx\]

OpenStudy (amistre64):

now this is one of those trig integrations from way back when

OpenStudy (amistre64):

is that french or italian?

OpenStudy (alfie):

That's italian.

OpenStudy (amistre64):

wolfram does simplifications, so im sure at one point that are the same

OpenStudy (alfie):

It just says "Nevertheless the discriminant is > 4" and then it says "Therefore". Nothing too much important. I see, but what's that "sett"?

OpenStudy (amistre64):

sett is 7

OpenStudy (amistre64):

-7 tg h yada yada

OpenStudy (amistre64):

maybe lol ... but i see that its not ....lol

OpenStudy (alfie):

so it's kind of a printing error or something? Why in the hell they'd write "sett"? More over in italian it's "sette" and ... I don't know :D . LoL.

OpenStudy (amistre64):

yep tanh^(-1) (2-x) and the book took out the negative and stuck it in front

OpenStudy (alfie):

Ohhhh... I just got what "sett" means. Sector tangent x. Which is the inverse function of the iperbolic tangent

OpenStudy (alfie):

Got it, all righty, some weird steps it does by the way, I'd have used lalaly method and I guess that'd have been simpler. Thank you all guys, that's really appreciated.

OpenStudy (lalaly):

Goodluck:D

OpenStudy (mr.math):

I don't think it's weird, this method can be used in situations where finding constants of the partial fractions is difficult. I came across such a problem last week and used the same approach the book used, (i.e make the top the derivative of the bottom so you get a log for the integration).

OpenStudy (mr.math):

And good luck! :)